OpenStudy (anonymous):

A man invests a certain amount of money at 2% interest and $800 more than that amount in another account at 4% interest. At the end of one year, he earned $92 in interest. How much money was invested in each account?

OpenStudy (anonymous):

high school math or calculus ? o_O

OpenStudy (anonymous):

Algebra 1 lol

OpenStudy (anonymous):

02x+.04*2x=112 .02x+.08x=112 .10x=112 x=112/.10 x=$1,120 invested @ 2%. 2*1,120=$2,240 amount invested @ 4%. Proof: .02*1,120+.04*2,240=112 22.4+89.6=112 112=112

OpenStudy (anonymous):

cool 'cuz every1 hates compound interest problem lmao

OpenStudy (anonymous):

Let interest earned of 1st acc be I(1)=P(1)*R(1)*T Let interest earned of 2nd acc be I(2)=P(2)*R(2)*T where P(1)=x , P(2)=x+800 , R(1)=0.02, R(2)=0.04 I(1)+I(2)=92 0.02x+0.04(x+800)= 92 0.02x+0.04x+32=92 0.06x=60 x=1000=P(1)=amount invested in 1st acc P(2)=x+800=1000+800=1800 = amount invested in 2nd acc